# Keystroke Dynamics: Concepts, Techniques, and Applications

### v2

@@ 2.4.2 Equal Error Rate
Figure 3: (a) Equal Error Rate (EER) as the instance of False Accept Rate (FAR) and False Reject Rate (FRR) intersecting, and the effect of loose and strict threshold (when the threshold is loose, the FAR is high and FRR is low; when the threshold is strict, the FAR is low and FRR is high). The EER balances between loose and strict threshold. (b) Receiver Operating Characteristic (ROC) curve and the Area Under the Curve (AUC) showing the performance of different classifiers. A classifier touching the point (0.0, 1.0) is desired and considered a perfect classifier. This is difficult in practical systems, a classifier closer to the coordinate (0.0, 1.0) is preferred. The ROC curve below the diagonal line is considered worse while the ROC curve above the diagonal line is considered better.
@@ 2.4.3 ROC and AUC
The Receiver Operating Characteristic (ROC) curve, is a graphical plot of FAR and TPR (i.e., 1 - FRR) that visualizes the performance of a binary classifier while the threshold for classification is changed. This metric provides a clear visualization of keystroke dynamics performance across varying classification thresholds. A perfect classifier will have a point at the top left corner with coordinate (0.0, 1.0) which illustrates no false accept or false reject (Figure 3(b)). Since this is hardly practical in a large scale system, a classifier closer to the coordinate (0.0, 1.0) is therefore desired. The diagonal from the top right to bottom left corners indicates a random classifier, which is similar to a random guess and therefore not desired. A classifier with an ROC curve lower than the diagonal line is mostly assumed to have been wrongly labeled and therefore can be flipped. The Area Under the Curve (AUC), often associated with the ROC curve, provides an combined performance measurement across all varying thresholds of classification.
@@ 2.4.4 ANIA and ANGA
In behavioral biometric studies, especially in continuous user authentication, performance metrics such as FAR, FRR, ROC and AUC may not be sufficient, as they fail to illustrate the average keystroke count required for a genuine user to be rejected or an impostor to be detected. Therefore, to further evaluate the performance of a CUA system, the Average Number of Impostor Actions (ANIA) and Average Number of Genuine Actions (ANGA) are used, where the actions in this case are keystrokes [84]. The ANGA is the number of legitimate actions users can perform before they are wrongfully denied access, while ANIA is the number of actions impostors can do before they are detected and locked-out. ANGA is perfered to be high (in fact, it is desired that a legitimate user should never be denied access), and ANIA as low as possible, so as to reduce the amount of actions an illegitimate user can perform before detection.
The ANIA and ANGA are mostly used in keystrokes dynamics continuous authentication and can be calculated as shown in Equation 3 and Equation 4, where $N$ is the amount of consecutive actions a user (genuine or impostor) can finish before an authentication/identification decision is made.
@@ 2.4.5 Authentication Time
The authentication time is an additional metric used in keystroke dynamics to describe the usability in the system. The metric is the total average time necessary for collecting keystroke data and the average of time system takes to permit or deny user access [123]. While the desire to achieve a very low FAR and FRR stand, it is important to achieve this with a very short authentication time. For example, the authentication system of Zheng et al. [130] achieved FAR and FRR of 1.30% but with a very long authentication time (approximately 38 minutes). Such a long authentication time would leave impostors uncaught as most impostors require signifigantly less time to complete attacks. Additionally, the mouse dynamics-based authentication of Shen et al. [103] had 3.33% FAR and 2.12% FRR with about 119 seconds authentication time. Although this FAR and FRR is higher than that of Zhen et al., it could be considered a better system based on the lower authentication time.
@@ 2.4.6 Rank-N
Rank-n [1] is an evaluation method used to measure the performance of keystroke dynamics based authentication systems. A dataset of multiple samples from each subject is used to test the system. For each subject, one of the samples is used as reference sample and the rest of the samples are considered as test samples. A similarity metric such as Euclidean distance is used to measure the similarity of a test sample to the reference sample. The test samples are ranked based on their similarity to the reference sample. The percentage of the inclusion of the correct test sample within the top n ranked samples is measured by the rank-n metric. If the correct test sample is included within the top 5 ranked samples 90% of the time, then the rank-5 accuracy is 90%. Therefore, the rank-n value needs to be lower for better performance. Because for lower rank-n value, there is a better chance that the correct test sample will be ranked higher. A rank-1 accuracy of 90% indicates that the correct test sample is ranked 1st 90% of the time. A comprehensive analysis of the system’s performance is provided by rank-n evaluation.
@@ 3 Keystroke Datasets
A keystroke dataset is required to perform satisfactory evaluation of keystroke dynamics-based authentication algorithms The quality, size and nature (fixed or free-text) of the dataset is a crucial factor in evaluating the performance of these algorithms. To perform evaluation, an algorithm is applied that measures and compares the difference in timing between key presses and releases in the dataset, using this information to make an authentication decision. This section details various keystroke datasets and their effectiveness in terms of authentication performance.
3.1 CMU Dataset (2009)
Killourhy and Maxion [66] collected fixed text (password) keystroke data and evaluated 14 algorithms with that dataset. 51 subjects were included in data collection. The fixed password was “.tie5Roanl”. Each subject typed this password 400 times during 8 sessions with 50 attempts in each session. There was at least one day gap between two sessions to reflect the temporal variation of typing of the subjects. A Microsoft Windows application was used for the data collection. For more accurate timestamps, an external reference clock was utilized. Killourhy and Maxion evaluated different algorithms with this dataset and achieved EER between 9.6% and 10.2%.
3.2 GreyC-A Dataset (2009)
Giot, El-Abed, and Rosenberger [45] proposed a benchmark database and software for measuring the performance of authentication systems leveraging keystroke dynamics for user access control. Their focus was on static-keystroke authentication (such as passwords or pass phrases). GREYC-Keystroke was the proposed software. They developed a standard keystroke dynamics database using the GREYC-Keystroke software. The data consisted of 133 participants including researchers, students, and employees. Each participant typed the password between 5 and 107 times. The authors collected the data of 7555 attempts. Most of the participants attended minimum 5 sessions. On average there were 51 attempts per participant. In [46], they used the GREYC keystroke dataset to evaluate their proposed technique based on SVM learning. The enrollment step was limited to 5 captures. The administrator set a passphrase that the users had to type. A two-class SVM was utilized for the enrollment. This SVM-based machine learning approach only needed five captured samples to develop the model. They used four types of algorithms for their experiment - 1) Statistical algorithm, 2) Distance-based algorithm, 3) Rhythm-based algorithm, 4) Machine learning based algorithm. They also investigated the variations in the authentication method due to different keyboards. The authors also considered the consequences of some other factors like differing amounts of samples needed to develop the model, individual or global threshold, the database size etc. The EER was calculated by using ten vectors from the head of the dataset for enrollment and the remaining samples used for the verification. In terms of EER, their proposed method performed better than all other algorithms.
3.3 GreyC-B Dataset (2012)
Giot, El-Abed, and Rosenberger [47] created a new dataset for keystroke based biometric systems. To develop the dataset, two types of logins and passwords were used- chosen and imposed. Leveraging a web environment, the data was collected without any restriction. They performed statistical analysis of some key factors including the effect of password size on performance, significance of fusion techniques etc. The data collection process consisted of three steps. 1) Ten attempts using a prior researcher chosen login and password, 2) Ten attempts for the user chosen login and password, 3) Ten imposter samples from two other participants. 83 individuals participated in the data collection. They collected 5185 genuine samples, 5439 imposed samples and 5754 imposter samples. For each participant, 20 samples were used for training and the remaining samples were used for testing (with a minimum of 20). The conclusions of their experiment were - 1) Using an individual threshold was better than using a global threshold for computing the EER, as an individual threshold is a value specific to a user that determines the origin of the sample - legitimate user or an imposter, 2) Using logins provided better results than using passwords, 3) The performance improved by doing the fusion of all features (monographs and digraphs), 4) The performance was dependent on the password size and entropy.
3.4 Pace Dataset (2013)
Bakelman et al. [18] collected fixed-text numeric keypad data for keystroke authentication of 30 subjects over for four days. On each day a maximum of 60 samples were collected for each subject. There were 11 keystrokes per sample from numeric sequence, and the Enter key. The data was collected by a third party keylogger. The subjects had to type the samples with only right hand in order to simulate typing a phone number or ATM pin. There were 20 samples per subject. For this dataset, they got the EER 10.5% and 6.1%.
3.5 Clarkon I Dataset (2014)
Vural et al. [122] developed a new dataset containing short pass-phrases, long text transcription(fixed-text) and free text. They also recorded videos showing subjects during data collection in order to capture the facial expressions and hand movements of the subjects. Vural et al. started the data collection at Clarkson University in 2011. The dataset includes 2 kinds of fixed text (password and transcriptions) and free text from 39 participants incliding university employess and students. For each participant, the data was collected in two one-hour sessions each occuring on a different day. The first task was typing 3 different passwords. The subjects typed each password 20 times. The passwords were: yesnomaybe, bahaNe312!, ballzonecart. For the second task, the subjects had to answer 9 survey questions in free text with a minimum of 500 characters to answer each survey question. The last task was copying the commencemen speech of Steve Jobs at Stanford University. A browser-based keylogger was used for the data collection. The data was collected for a period of 11 months between August 2011 and June 2012 in a lab at Clarkson University. On average each subject provided 11066 keystrokes in session 1 and 10467 keystrokes in session 2. Every dession recording was analyzed to determine digraphs and trigraphs. N-graphs greater than 500 milliseconds or shorter than 30 milliseconds in length were discarded. In session 1, on average there were 8406 digraphs and 389 unique digraphs per subject, 7561 trigraphs and 1603 unique trigraphs per subject. In session 2, on average there were 8261 digraphs and 388 unique digraphs per subject, 7520 trigraphs and 1620 unique trigraphs per subject. To evaluate this dataset, two algorithms were used: Leggett et al. [72] and Gunetti & Picardi [52]. Using the algorithm of Leggett et al. [72], an FAR of 0.25% and FRR of 17.65% were obtained for a threshold of 0.9; an FAR of 3.45% and FRR of 8.82% were obtained for a threshold of 0.85. Ceker and Upandhyaya [24] utilized this dataset to examine the accuracy of a single-class support vector machine (SVM) on longer text data, as opposed to shorter password text. They found that using the four ranking digraphs by comonality (‘he’, ‘re’, ‘th’ and ‘an’) resulted in a 2.94% EER, while using the most common 12 or more digraphs led to an almost 0% EER.
3.6 OhKBIC Dataset (2015)
Monaco et al. [82] described the results of the One-handed Keystroke Biometric Identification Competition (OhKBIC). In the dataset, Monaco et al. included the data of normal typing and handicapped typing (typing with one hand). Data was collected from three online exams of undergraduate students. There were 64 subjects. In each exam, there were five essay questions. The subjects typed normally in the first exam. The subjects were directed to type with their left hand only for the second exam and with their right hand only for the third exam. Each subject typed a minimum 500 keystrokes on each exam. The data was collected by a JavaScript based keylogger and transferred to a server. Each used different keyboard make and models throughout the three exams which was a drawback of this dataset. A portion of the data collected from the first exam of normal both-hands typing was used as labeled training data. The unlabeled data included all three types of data (both-hand) typing, left-hand typing and right-hand typing data. The top team used two regression models (Artificial Neural Network and Counter-Propagation Artificial Neural Network) and one prediction model (Support Vector Machine) for pairwise coupling [54]. The best recognition rate ranged from 55.7% to 82.8% for both-hand section, 19.9% to 30.5% for left-hand section and 24.8% to 34.3% for right-hand section.
3.7 Buffalo Dataset (2016)
Sun, Ceker, and Upadhyaya [114] proposed a shared dataset containing keystrokes collected from 157 participants. The data was both fixed and free-text. The data collection process continued for 4 months between September and December 2015. For each subject, the data was colled in 3 sessions. The tasks included fixed-text transcription and answering quesions in free text. On average there was a 28 days gap between 2 sessions to consider the temporal variations. There are 2 sections in the dataset- in one section the subjects used same keyboard across sessions. In the other section, the subjects used different keyboards across sessions. The first task was copying the 2005 commencement speech of Steve Jobs at Stanford University. The speech was divided into 3 parts with each part of equal length, the subjects had to type one part in each session. The second task consisted of multiple subtasks including answering survey questions and describing a picture, sending an email with attachment and free internet browsing. Each session lasted for about 50 minutes with 30 minutes for the first task and 20 minutes for the second task. A system logger was adopted for the data collection on the windows platform [41][115]. Among the 157 subjects, the data of 148 subjects was included in the dataset. The data of the remaining users was discarded due to mistakes. On average each subject typed around 5700 keystrokes in each session, around 17000 keystrokes in 3 sessions. There was a 3 to 5 week time gap between 2 sessions to reflet the temporal effect. Gaussian mixture model, a probabilistic model for classifying data into different categories based on the probability distribution, was utilized to evaluate this dataset. The EER was 0.01% and 0.39% for Gaussian mixture model with one and two components, respectively.
3.8 Clarkson II Dataset (2017)
A novel dataset was proposed by Murphy et al. [85] that included keystrokes, mouse events and active programs. 103 users contributed the data for a period of around 2.5 years. Unlike the other datasets, the data was collected on the users’ personal computers during normal use. Keystrokes and mouse data were recorded while users make use of their computers, such as for typing, gaming etc thereby classifying the data as free-text. The users provided 12.9 M keystrokes with each used providing an average of 125K keystrokes. The keylogger was installed on each user’s computer to recorded keystroke and mouse movement movement events. The recorded data was sent to a remote database server. 10 samples, each sample consisting of 1000 keystrokes, were used to form the reference profile of a user. The rest of the data of a user was considered genuine test samples. For each user, the data of all other users was considered as imposter data. The EER was 10.36% for this dataset. This result was worse compared to fixed-text and controlled free-text datasets because it is a completely uncontrolled dataset and as such contains several keystrokes that are not usually seen in typing tasks.
@@ 3.8 Clarkson II Dataset (2017)
Table 2: A summary of major keystroke datasets by names and references of these datasets, “the type of data” like free text, fixed text or semi fixed text, number of subjects that participated in the data collection and “average number of keystrokes” that each subject provided.
@@
Pace (2013) by Bakelman et al. [18] Fixed 30 220 Password and numeric input Not available
@@
OhKBIC (2015) by Monaco et al. [82] Free 64 minimum 1,500 One-handed Not available
@@
AR (2021) by Wahab et al. [125] Semi-Fixed 42 2,048 Account recovery Not available
@@
Multi-K (2022) by Wahab et al. [124] Free 86 24,000 Multi-keyboard and bilingual Not available

### v3

@@ 2.4.2 Equal Error Rate
Figure 3: (a) Equal Error Rate (EER) as the instance of False Accept Rate (FAR) and False Reject Rate (FRR) intersecting, and the effect of loose and strict threshold (when the threshold is loose, the FAR is high and FRR is low; when the threshold is strict, the FAR is low and FRR is high). The EER balances between loose and strict thresholds. (b) The Receiver Operating Characteristic (ROC) curve and the Area Under the Curve (AUC) show the performance of different classifiers. A classifier touching the point (0.0, 1.0) is desired and considered a perfect classifier. This is difficult in practical systems; a classifier closer to the coordinate (0.0, 1.0) is preferred. The ROC curve below the diagonal line is considered worse, while the ROC curve above the diagonal line is considered better.
@@ 2.4.3 ROC and AUC
The Receiver Operating Characteristic (ROC) curve is a graphical plot of FAR and TPR (i.e., 1 - FRR) that visualizes the performance of a binary classifier while the threshold for classification is changed. This metric provides a clear visualization of keystroke dynamics performance across varying classification thresholds. A perfect classifier will have a point at the top left corner with coordinates (0.0, 1.0) which illustrates no false accept or false reject (Figure 3(b)). Since this is hardly practical in a large-scale system, a classifier closer to the coordinate (0.0, 1.0) is desired. The diagonal from the top right to bottom left corners indicates a random classifier, which is similar to a random guess and therefore not desired. The Area Under the Curve (AUC), often associated with the ROC curve, provides a combined performance measurement across all varying thresholds of classification.
@@ 2.4.4 ANIA and ANGA
In keystroke dynamics, especially in continuous user authentication, performance metrics such as FAR, FRR, ROC, and AUC may not be sufficient, as they fail to illustrate the average keystroke count required for a genuine user to be rejected or an impostor to be detected. Therefore, to further evaluate the performance of a continuous user authentication (CUA) system, the Average Number of Impostor Actions (ANIA) and Average Number of Genuine Actions (ANGA) are used, where the actions in this case are keystrokes [84]. The ANGA is the number of legitimate actions users can perform before they are wrongfully denied access, while ANIA is the number of actions impostors can perform before they are detected and locked out. ANGA is preferred to be high (in fact, it is desired that a legitimate user should never be denied access), and ANIA as low as possible, so as to reduce the amount of actions an illegitimate user can perform before detection.
The ANIA and ANGA are mostly used in keystroke dynamics continuous authentication and can be calculated as shown in Equation 3 and Equation 4, where $N$ is the number of consecutive actions a user (genuine or impostor) can finish before an authentication/identification decision is made.
@@ 2.4.5 Authentication Time
The authentication time is an additional metric used in keystroke dynamics to describe the usability of the system. This metric is the total average time necessary for collecting keystroke data and the average of the time a system takes to permit or deny user access [123]. While the desire to achieve a very low FAR and FRR stands, it is important to achieve this with a very short authentication time. For example, the authentication system of Zheng et al. [130] achieved FAR and FRR of 1.30% but with a very long authentication time (approximately 38 minutes). Such a long authentication time would leave impostors uncaught, as most impostors require significantly less time to complete attacks. Additionally, the mouse dynamics-based authentication of Shen et al. [103] had 3.33% FAR and 2.12% FRR with about 119 seconds of authentication time. Although these FAR and FRR are higher than those of Zhen et al., it could be considered a better system based on the lower authentication time.
@@ 2.4.6 Rank-N
Rank-n [1] is an evaluation method used to measure the performance of keystroke dynamics-based authentication systems. A dataset of multiple samples from each subject is used to test the system. For each subject, one of the samples is used as a reference sample and the rest of the samples are considered as test samples. A similarity metric such as Euclidean distance is used to measure the similarity of a test sample to the reference sample. The test samples are ranked based on their similarity to the reference sample. The percentage of the inclusion of the correct test sample within the top n-ranked samples is measured by the rank-n metric. If the correct test sample is included within the top 5 ranked samples 90% of the time, then the rank-5 accuracy is 90%. Therefore, the rank-n value needs to be lower for better performance. For a lower rank-n value, there is a better chance that the correct test sample will be ranked higher. A rank-1 accuracy of 90% indicates that the correct test sample is ranked 1st 90% of the time. Rank-n evaluation provides a comprehensive analysis of the system’s performance.
@@ 3 Keystroke Datasets
A keystroke dataset is required to perform a satisfactory evaluation of keystroke dynamics-based authentication algorithms. The quality, size, and nature (fixed, free, or semi-fixed text) of the dataset are crucial factors in evaluating the performance of these algorithms. To perform an evaluation, an algorithm is applied that compares the typing patterns of users based on keystroke features (see Fig. 1). This information is used to authenticate users. This section details various keystroke datasets and their effectiveness in terms of authentication performance. A summary of the keystroke datatsets is shown in Table 2.
@@
Table 2: A summary of publicly available keystroke datasets. The table includes the type of data collected, the number of subjects that participated in the data collection, and the average number of keystrokes per subject.
@@
Pace (2013) by Bakelman et al. [18] Fixed 30 220 Password and numeric input Unknown
@@
OhKBIC (2015) by Monaco et al. [82] Free 64 minimum 1,500 One-handed Unknown
@@
AR (2021) by Wahab et al. [125] Semi-Fixed 44 2,048 Account recovery Dataset available upon request from the authors
@@
Multi-K (2022) by Wahab et al. [124] Free 86 24,000 Multi-keyboard and bilingual Dataset available upon request from the authors
3.1 CMU Dataset (2009)
Killourhy and Maxion [66] collected fixed text (password) keystroke data and evaluated 14 algorithms with that dataset. 51 subjects were included in the data collection. The fixed password was “.tie5Roanl”. Each subject typed this password 400 times during 8 sessions, with 50 attempts in each session. There was at least a one-day gap between successive sessions to reflect the temporal variation in typing of the subjects. A Microsoft Windows application was used for the data collection. An external reference clock was utilized to make more accurate timestamps. Killourhy and Maxion evaluated different algorithms with this dataset and achieved an EER between 9.6% and 10.2%.
3.2 GreyC-A Dataset (2009)
Giot, El-Abed, and Rosenberger [45] proposed a benchmark database and software for measuring the performance of authentication systems leveraging keystroke dynamics for user access control. Their focus was on static keystroke authentication (such as passwords or passphrases). GREYC-Keystroke was the proposed software. They developed a standard keystroke dynamics database using the GREYC-Keystroke software. Keystroke data was collected for 133 participants including researchers, students, and employees. Each participant typed the password between 5 and 107 times. The authors collected the data of 7555 attempts. Most participants attended at least 5 sessions. On average, there were 51 attempts per participant.
Giot et al. [46] used the GREYC keystroke dataset to evaluate their proposed technique based on SVM learning. The enrollment step was limited to 5 captures. The administrator set a passphrase that the users had to type. A two-class SVM was utilized for the enrollment. This SVM-based machine learning approach only needed five captured samples to develop the model. They used four types of algorithms for their experiment: 1) Statistical algorithm, 2) Distance-based algorithm, 3) Rhythm-based algorithm, and 4) Machine learning based algorithm. They also investigated the variations in the authentication method due to different keyboards. The authors also considered the consequences of some other factors like differing amounts of samples needed to develop the model, individual or global threshold, the database size, etc. The EER was calculated by using ten vectors from the head of the dataset for enrollment and the remaining samples were used for verification. In terms of EER, their proposed method performed better than all other algorithms.
3.3 GreyC-B Dataset (2012)
Giot, El-Abed, and Rosenberger [47] created a new dataset for keystroke-based biometric systems. To build the dataset, two types of logins and passwords were chosen and imposed. Leveraging a web environment, the data was collected without any restrictions. They performed statistical analysis of some key factors including the effect of password size on performance, the significance of fusion techniques, etc. The data collection process consisted of three steps. 1) Ten attempts using a prior researcher-chosen login and password, 2) Ten attempts for the user-chosen login and password, and 3) Ten imposter samples from two other participants. 83 individuals participated in the data collection. They collected 5185 genuine samples, 5439 imposed samples, and 5754 imposter samples. For each participant, 20 samples were used for training and the remaining samples were used for testing (with a minimum of 20). The conclusions of their experiment were: 1) Using an individual threshold was better than using a global threshold for computing the EER because an individual threshold is a value specific to a user that takes into account the origin of the sample i.e whether a legitimate user or an imposter, 2) Using logins provided better results than using passwords, 3) The performance improved by doing the fusion of all features (monographs and digraphs), 4) The performance was dependent on the password size and entropy.
3.4 Pace Dataset (2013)
Bakelman et al. [18] collected fixed-text numeric keypad data for keystroke authentication of 30 subjects over four days. On each day a maximum of 60 samples were collected from each subject. There were 11 keystrokes per sample from numeric sequence and the Enter key. The data was collected by a third-party keylogger. The subjects had to type the samples with only their right hand in order to simulate typing a phone number or ATM PIN. There were 20 samples per subject. For this dataset, they got an EER of 6.1%.
3.5 Clarkson I Dataset (2014)
Vural et al. [122] developed a new dataset containing short pass-phrases, long text transcription (fixed-text) and free text. They also recorded videos showing subjects during data collection in order to capture the facial expressions and hand movements of the subjects. The dataset includes 2 kinds of fixed text (password and transcriptions) and free-text from 39 participants including university employees and students. For each participant, the data was collected in two one-hour sessions each occurring on a different day. The first task was typing 3 different passwords. The subjects typed each password 20 times.
For the second task, the subjects had to answer 9 survey questions in free-text with a minimum of 500 characters for each answer. The last task was copying the commencement speech of Steve Jobs at Stanford University. A browser-based keylogger was used for the data collection. The data was collected for a period of 11 months between August 2011 and June 2012 in a lab at Clarkson University. On average, each subject provided 11066 keystrokes in session 1 and 10467 keystrokes in session 2. Every session recording was analyzed to determine digraphs and trigraphs. N-graphs greater than 500 milliseconds or shorter than 30 milliseconds in length were discarded. In session 1, on average there were 8406 digraphs and 389 unique digraphs per subject, 7561 trigraphs, and 1603 unique trigraphs per subject. In session 2, on average there were 8261 digraphs and 388 unique digraphs per subject, 7520 trigraphs, and 1620 unique trigraphs per subject.
To evaluate this dataset, the algorithm of Leggett et al. [72] was used. A FAR of 0.25% and FRR of 17.65% were obtained for a threshold of 0.9; aFAR of 3.45% and FRR of 8.82% were obtained for a threshold of 0.85. Ceker and Upandhyaya [24] utilized this dataset to examine the accuracy of a single-class support vector machine (SVM) on longer text data, as opposed to shorter password text. They found that using the four ranking digraphs by commonality (‘he’, ‘re’, ‘th’ and ‘an’) resulted in a 2.94% EER while using the most common 12 or more digraphs led to an almost 0% EER.
3.6 OhKBIC Dataset (2015)
Monaco et al. [82] described the results of the One-handed Keystroke Biometric Identification Competition (OhKBIC). In the dataset, Monaco et al. included the data of normal typing and handicapped typing (typing with one hand). Data was collected from three online exams of undergraduate students. There were 64 subjects. In each exam, there were five essay questions. The subjects typed normally in the first exam. The subjects were directed to type with their left hand only for the second exam and with their right hand only for the third exam. Each subject typed a minimum of 500 keystrokes on each exam. The data was collected by a JavaScript-based keylogger and transferred to a server. Each used different keyboard make and models throughout the three exams which was a drawback of this dataset. A portion of the data collected from the first exam of normal both-hands typing was used as labeled training data. The unlabeled data included all three types of data (both-hand) typing, left-hand typing, and right-hand typing data. The top team used two regression models (Artificial Neural Network and Counter-Propagation Artificial Neural Network) and one prediction model (Support Vector Machine) for pairwise coupling [54]. The best recognition rate ranged from 55.7% to 82.8% for both-hand sections, 19.9% to 30.5% for the left-hand section, and 24.8% to 34.3% for the right-hand section.
3.7 Buffalo Dataset (2016)
Sun, Ceker, and Upadhyaya [114] developed a shared dataset containing keystrokes collected from 157 participants. The data was both fixed and free-text. The data collection process continued for 4 months between September and December 2015. For each subject, the data was collected in 3 sessions. The tasks included fixed-text transcription and answering questions in free text. On average there was a 28 days gap between 2 sessions to consider the temporal variations. There are 2 sections in the dataset- in one section the subjects used the same keyboard across sessions. In the other section, the subjects used different keyboards across sessions. The first task was copying the 2005 commencement speech of Steve Jobs at Stanford University. The speech was divided into 3 parts of equal length; the subjects had to type one part in each session. The second task consisted of multiple subtasks including answering survey questions and describing a picture, sending an email with an attachment, and free internet browsing. Each session lasted for about 50 minutes with 30 minutes for the first task and 20 minutes for the second task. A system logger was adopted for the data collection on the Windows platform [41][115]. Among the 157 subjects, the data of 148 subjects was included in the dataset. The data of the remaining users was discarded due to mistakes. On average each subject typed about 5700 keystrokes in each session, around 17000 keystrokes for all 3 sessions. There was a 3 to 5 week time gap between successive sessions to reflect the temporal effect. The Gaussian mixture model, a probabilistic model for classifying data into different categories based on the probability distribution, was utilized to evaluate this dataset. The EER was 0.01% and 0.39% for the Gaussian mixture model with one and two components, respectively.
3.8 Clarkson II Dataset (2017)
A novel dataset was proposed by Murphy et al. [85] that included keystrokes, mouse events, and active programs. 103 users contributed the data for a period of around 2.5 years. Unlike the other datasets, the data was collected on the users’ personal computers during normal use. Keystrokes and mouse data were recorded while users make use of their computers, such as for typing, gaming, etc thereby classifying the data as free-text. The users provided 12.9 M keystrokes, with each user providing an average of 125K keystrokes. A keylogger was installed on each user’s computer to record keystroke and mouse movement events. The recorded data was sent to a remote database server. 10 samples, each sample consisting of 1000 keystrokes, were used to form the reference profile of a user. The rest of the user’s data was considered genuine test samples. For each user, the data of all other users was considered as imposter data. The EER was 10.36% for this dataset. This result was worse compared to fixed-text and controlled free-text datasets because it is a completely uncontrolled dataset and as such contains several keystrokes that are not usually seen in typing tasks.